OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6.

OpenStudy (matt101):

Point-slope form in general is: y – y1 = m(x – x1) We have m, x1, and y1, so just plug them in: y + 1 = 6(x - 5)

OpenStudy (matt101):

You can rearrange and solve for y to get the slope-y intercept form, which I think is far more useful...